OpenStudy (anonymous):

Which method is the best method for solving the equation 8x² – 13x + 3 = 0? square roots factoring graphing quadratic formula

OpenStudy (agreene):

E. Completing the Square. (add 8 to both sides and factor)

OpenStudy (agreene):

erg, divide both sides by 8 and factor... no matter, they probably want you to say Quadratic Formula.

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

i dont think it is factorable, two number that add to be 13 but multiply to be 3 :?

hero (hero):

graphing calc

OpenStudy (anonymous):

factors like this -1/32 (13+sqrt(73)-16 x) (-13+sqrt(73)+16 x)

OpenStudy (agreene):

8x^2-13x+3 x^2-13x/8+3/8=0 x^2-13x/8=3/8 x^2-13x/8+169/256=73/256 (x-13/16)^2=73/256 ...